# Bayesian Optimization for Robust State Preparation in Quantum Many-Body Systems

### v1

@@ Abstract
Introduction – Recent advances in quantum simulation platforms based on ultracold atoms in optical lattices Bloch et al. (2012); Gross and Bloch (2017) have allowed the creation of increasingly complex quantum many-body states in larger systems, necessitating improvements in state-preparation quality and speed. Well-established approaches like adiabatic ramping Eckstein and Kollar (2010), manual mapping of the quantum many-body gap Léonard et al. (2023), and optimal control techniques Doria et al. (2011); Glaser et al. (2015); Koch (2016); Koch et al. (2022) have allowed experimental access to exciting new phenomena, such as magnetic and topological states Simon et al. (2011); van Frank et al. (2016); de Léséleuc et al. (2019); Léonard et al. (2023).
@@
In recent years, machine learning methods, such as reinforcement learning Niu et al. (2019); Paparelle et al. (2020), have emerged as a powerful tool for quantum optimal control (QOC) applications. Bayesian optimization (BO) Shahriari et al. (2016); Garnett (2023), an alternative machine-learning approach, has been used to generate efficient protocols for fast creation of high-particle number Bose-Einstein condensates Wigley et al. (2016); Vendeiro et al. (2022), cooling in magneto-optical traps Tranter et al. (2018), and state preparation Mukherjee et al. (2020a); Xie et al. (2022). In these works, an online optimization loop is formed between the learner, and either an experimental setup or numerical simulation. The resulting optimized protocols have been demonstrated to compare favorably to manual ramp design Vendeiro et al. (2022) and other optimization techniques, such as differential evolution and Nelder-Mead Xie et al. (2022); Lagarias et al. (1998); Das and Suganthan (2011). BO is well suited to these tasks featuring experimental disorder or low repetition rates, as it typically requires fewer iterations, than e.g., methods based on gradient descent, and training is robust to moderate noise.
@@
In this work, we use an exact-diagonalization (ED) simulation and apply BO to prepare a few-body fractional quantum Hall (FQH) state which has recently been realized in an ultracold-atom experiment Léonard et al. (2023). This allows us to compare the optimized result to a protocol that is manually designed and has been implemented in practice. We extensively investigate how taking disorder in the experimental system into account changes the optimal strategy and how to engineer effective control paths in a noisy environment. Based on this analysis, we explore how exploiting classical simulation resources can boost efficiency by reducing experimental costs. From this, we motivate a general optimization strategy, which we believe to be applicable to any experimental system in which classical simulation is available in some capacity. These include not only the preparation of the Laughlin state Laughlin (1983), which acts as our benchmark, but also other current issues including challenging state preparation, such as the realization of quantum spin liquid states Broholm et al. (2020), and cooling in Fermi-Hubbard type systems Mazurenko et al. (2017); Chiu et al. (2018).
@@
In contrast to other optimal-control applications realizing a phase transition Xie et al. (2022), e.g., the superfluid to Mott-insulator transition Rosi et al. (2013); van Frank et al. (2016); Sørensen et al. (2019); Mukherjee et al. (2020a), the FQH problem features a topologically non-trivial target state prepared by tuning a larger number of control fields (4 vs 1-2) to be optimized over time. This more demanding problem can shed light on the prospects of machine-learning-optimized state preparation, entering a regime where simple linear ramping is impossible and an optimized protocol’s robustness and generality become primary concerns. To determine whether BO is a tool fit to tackle these challenges, our work proceeds as follows: We discuss our optimization strategy and investigate the experimental setting motivating our work by comparing experimental data to ED simulation. In the core part of our work, we apply BO to the simulated system and thoroughly investigate the performance of optimized ramping protocols regarding fidelity, speed, and robustness to disorder. We conclude with a proposal that we believe will facilitate the creation of more complex states in the system under investigation and also act as a general tool to boost efficiency in QOC applications going beyond this specific system.
Optimization Strategy – In this work, we implement a BO training loop as illustrated in Fig. 1.
@@
Figure 1: Optimization Strategy: (top) Illustration of the Bayesian optimization training loop. The experiment or numerical simulation acts as a black-box function, providing a quality measure for a given ramping protocol. Based on evaluations of this function, the BO learner builds a surrogate model of the cost landscape and suggests the next protocol to implement. (bottom) Proposed optimization strategy exploiting numerical resources to boost optimization efficiency in the experiment. Depending on the task at hand, several avenues with varying experimental costs may be explored. The most straightforward way is to directly implement the optimization result obtained from simulation in the experiment, requiring $\mathcal{O}(1)$ evaluations of the protocol’s quality. On top of this, the BO surrogate model can provide a sensitivity analysis that can be exploited to manually calibrate only the most sensitive protocol parameters in $\mathcal{O}(10)$ experimental evaluations. Finally, when substantial adaptation to the experimental setting is necessary, online optimization can be performed in $\mathcal{O}(100)$ iterations directly with the experiment. Transfer-learning approaches can leverage numerical resources to accelerate training significantly, e.g., by initializing the surrogate model used for training in the experiment with training data from the simulation. In this work, we highlight and motivate these three scenarios by comparing the numerical simulation of an idealized system to a more realistic description of the experimental setting that takes experimental disorder into account.
@@
Gaussian processes (GP) Rasmussen and Williams (2005) are used as a surrogate model of the cost function or figure of merit of ramping protocols over the entire parameter space. For details about BO and our implementation, the reader is referred to Appendix A.
@@
Experimental Setting – Our application of BO is focused on the state-preparation problem of a bosonic two-particle FQH state recently realized in experiment Léonard et al. (2023). The experiment features a ramp protocol, adiabatically  [^1] In the context of optimization, we consider an adiabatic protocol to be any strategy aiming to minimize losses of the target-state population by limiting the ramp speed to the maximum set by the inverse of the many-body gap. transforming an initial state of two pinned particles into the delocalized FQH state.
@@
Figure 2: Experimental Ramp: (a) Experimental ramping protocol for the $4$ lattice parameters adapted from Léonard et al. (2023). (b) Fidelity w.r.t. the instantaneous ground state throughout the ramp. The ED simulation is compared to experimental data Léonard et al. (2023). The color-coded lines in the background show the energy levels $E-E_{0}$ of $\hat{H}(t)$ and their respective occupation during the simulated ramping protocol. Here, $E_{0}$ is the instantaneous ground-state energy. Before $t_{x}$ is turned on at $T=60\>\tau$, the fidelity is defined in the 1D system formed by the 1st column of the $4\times 4$ lattice, instead of the full 2D system. This is done to avoid spurious drops to $F\sim 0$ when $\Delta_{x}<\Delta_{y}$ caused by the initial-state geometry.
@@
Figure 3: Optimized Ramp: (a) Optimized ramp protocol for the $4$ lattice parameters. White dots show the parameter values controlled by the optimizer at fixed points in time (vertical dashed lines). (b) Evaluated cost function at each iteration of the BO training. The visible oscillations in the evaluated fidelity follow sweeps of the BO hyperparameter $\kappa$, which tunes the strategy between exploration and exploitation (see Appendix A for optimization details). (c) Fidelity w.r.t. the instantaneous ground state throughout the ramp. The color-coded lines in the background show the energy levels of $\hat{H}(t)$ and their respective occupation during the simulated ramping protocol. All energies are w.r.t. the ground state energy which sets $E=0$. Before $t_{x}$ is turned on at $T=60\>\tau$, the fidelity is defined in the 1D system formed by the 1st column of the $4\times 4$ lattice, instead of the full 2D system. This is done to avoid spurious drops to $F\sim 0$ when $\Delta_{x}<\Delta_{y}$ caused by the initial-state geometry.
@@
To achieve this, four global lattice parameters in the interacting Harper-Hofstadter model Tai et al. (2017); Harper (1955); Hofstadter (1976)
$\begin{split}\hat{H}=\ &-t_{x}\sum_{x,y}(e^{i\phi y}\hat{a}^{\dagger}_{x,y}\hat{a}_{x+1,y}+\mathrm{H.c.})+\Delta_{x}\sum_{x,y}x\,\hat{n}_{x,y}\\ &-t_{y}\sum_{x,y}(\hat{a}^{\dagger}_{x,y}\hat{a}_{x,y+1}+\mathrm{H.c.})+\Delta_{y}\sum_{x,y}y\,\hat{n}_{x,y}\\ &+\frac{U}{2}\sum_{x,y}\hat{n}_{x,y}(\hat{n}_{x,y}-1)\end{split}$ (1)
@@
are varied independently over time: (i, ii) the tunneling strengths $t_{x,y}$ in x- and y-direction and (iii, iv) the chemical potential gradients $\Delta_{x,y}$. Here, $\hat{a}^{(\dagger)}_{x,y}$ is the bosonic annihilation (creation) operator at site $(x,y)$ and $\hat{n}_{x,y}=\hat{a}^{\dagger}_{x,y}\hat{a}_{x,y}$ is the particle number operator. The tunneling phase $\phi/2\pi=0.27$ is constant throughout this work, corresponding to the Laughlin-like state Laughlin (1983) at the fractional filling $\nu=1/2$, where $\nu$ is the filling factor. This target state is realized in the flat lattice $\Delta_{x}=\Delta_{y}=0$, at equal tunneling amplitudes $t_{x}=t_{y}=\hbar/\tau$. We fix the value of the tunneling amplitudes in the target state as our unit of energy while the tunneling time $\tau$ acts as the unit of time. In the experiment, $\tau=4.3(1)\>\mathrm{ms}$ Léonard et al. (2023). Finally, the on-site repulsion is fixed to $U=8\>\hbar/\tau$.
@@
The ramping protocol used in the experiment was devised by mapping out the many-body gap over the four-dimensional parameter space and manually identifying a path that avoids a gap closing. In addition, the tunneling $t_{x}$ in x-direction is only turned on in the latter half of the experimental protocol to reduce the effects of Floquet heating Sträter and Eckardt (2016) resulting from the synthetic magnetic field engineered in the experiment. The protocol is illustrated in Fig. 2. It serves as the primary benchmark in this work – representing the capabilities of sophisticated adiabatic ramp design.
@@
where |$\ket{\psi_{\mathrm{final}}}$ is the state after the preparation protocol, and |$\ket{\psi_{0}}$ is the ground state of $\hat{H}$ at the target parameters $t_{x}=t_{y}=\hbar/\tau$ and $\Delta_{x}=\Delta_{y}=0$. This is a good figure of merit since it is experimentally accessible by inverting the ramp and measuring the fraction that has returned to the initial state. For the experimental protocol it was measured to be $F=43(6)\%$ Léonard et al. (2023). In Fig. 2, we compare the experimental data to an ED time-evolution simulation. Even though the simulation is performed at zero temperature, does not take into account any lattice defects, and does not model the Floquet sequence employed in the experiment – instead simulating the unperturbed effective Hamiltonian – the fidelity-loss over time (see Fig. 2b)) shows qualitative agreement with the experiment. The fidelity of $F=53.5\%$ observed in the simulation is not much higher than the value reported in the experiment. This supports non-adiabatic losses to excited states during the state-preparation protocol as the leading source of fidelity loss, providing a promising foundation for ramp optimization.
Optimization – For each lattice parameter, we parametrize the time-dependent ramp by $N_{s}$ linear segments of equal length. Fixing the initial and target parameters to be the same as in the experiment leads to a BO problem with $4\times(N_{s}-1)$ parameters. Motivated by the experimental conditions and concerns about Floquet heating, we limit the search space to
@@
which corresponds to a Floquet regime that minimizes excitations to higher bands and achieves efficient dynamics. Specifically, the range for $t_{x/y}$ is chosen to limit the broadening of heating channels Sträter and Eckardt (2016), as its strength is determined by the driving amplitude. We start ramping $t_{x}$ and $\Delta_{x}$ only after passing the time $T_{x}=0.6\>T$ where $T$ is the total ramp time. After training with $500$ iterations, BO identifies an optimal ramp with $F=94.5\%$ as illustrated in Fig. 3. Interestingly, the protocol is qualitatively similar to the one used in the experiment. As depicted in Fig. 3, it achieves near-perfect fidelity due to significant gains in the ground-state population (w.r.t. the instantaneous ground state) near the final protocol time. Such a resurgence of fidelity has also been reported in other simulation works on QOC applications Mukherjee et al. (2020b). While captured by optimal control theory, it is outside the scope of manual design principles, such as adiabatic ramp design, which aim to minimize fidelity losses, but do not take processes that can return population to the ground state into account. We now turn to whether this feature, observable in simulations, can be expected to be robust in experimental conditions.
Training with Noise – To characterize robustness, and estimate the real-world fidelity we would expect from optimized protocols, we now model experimental disorder in our simulation by the sum of random and harmonic chemical potential offsets (see Appendix C). The disorder $\mu$ is taken to be Gaussian-distributed with mean 00 and standard deviation $\sigma=0.1\>\hbar/\tau$ which is an experimentally realistic disorder level. For each numerical evaluation of a ramp’s fidelity, we perform time evolution with a newly sampled realization of the disorder.
Optimization is now performed in the presence of this disorder to represent training in an experimental setting. To mitigate fluctuations, we take the average fidelity of $5$ disorder realizations as the cost function, which is a typical value used in the literature Vendeiro et al. (2022); Xie et al. (2022). After training for $1000$ iterations, we compare the optimized protocol to the manually designed ramp and the optimum found from training without disorder in Fig. 4.

### v2

@@ 1 Introduction
Recent advances in quantum simulation platforms based on ultracold atoms in optical lattices [1, 2] have allowed the creation of increasingly complex quantum many-body states in larger systems, necessitating improvements in state-preparation quality and speed. Well-established approaches like adiabatic ramping [3], manual mapping of the quantum many-body gap [4], and optimal control techniques [5, 6, 7, 8] have allowed experimental access to exciting new phenomena, such as magnetic and topological states [9, 10, 11, 4].
@@
In recent years, machine learning methods, such as reinforcement learning [12, 13], have emerged as a powerful tool for quantum optimal control (QOC) applications. Bayesian optimization (BO) [14, 15], an alternative machine-learning approach, has been used to generate efficient protocols for fast creation of high-particle number Bose-Einstein condensates [16, 17], cooling in magneto-optical traps [18], and state preparation [19, 20]. In these works, an online optimization loop is formed between the learner, and either an experimental setup or numerical simulation. The resulting optimized protocols have been demonstrated to compare favorably to manual ramp design [17] and other optimization techniques, such as differential evolution and Nelder-Mead [20, 21, 22]. BO is well suited to these tasks featuring experimental disorder or low repetition rates, as it typically requires fewer iterations, than e.g., methods based on gradient descent, and training is robust to moderate noise.
@@
In this work, we use an exact-diagonalization (ED) simulation and apply BO to prepare a few-body fractional quantum Hall (FQH) state which has recently been realized in an ultracold-atom experiment [4]. This allows us to compare the optimized result to a protocol that is manually designed and has been implemented in practice. We extensively investigate how taking disorder in the experimental system into account changes the optimal strategy and how to engineer effective control paths in a noisy environment. Based on this analysis, we explore how exploiting classical simulation resources can boost efficiency by reducing experimental costs. From this, we motivate a general optimization strategy, which we believe to be applicable to any experimental system in which classical simulation is available in some capacity. These include not only the preparation of the Laughlin state [23], which acts as our benchmark, but also other current issues including challenging state preparation, such as the realization of quantum spin liquid states [24], and cooling in Fermi-Hubbard type systems [25, 26].
@@
In contrast to other optimal-control applications realizing a phase transition [20], e.g., the superfluid to Mott-insulator transition [27, 10, 28, 19], the FQH problem features a topologically non-trivial target state prepared by tuning a larger number of control fields (4 vs 1-2) to be optimized over time. This more demanding problem can shed light on the prospects of machine-learning-optimized state preparation, entering a regime where simple linear ramping is impossible and an optimized protocol’s robustness and generality become primary concerns. To determine whether BO is a tool fit to tackle these challenges, our work proceeds as follows: We discuss our optimization strategy and investigate the experimental setting motivating our work by comparing experimental data to ED simulation. In the core part of our work, we apply BO to the simulated system and thoroughly investigate the performance of optimized ramping protocols regarding fidelity, speed, and robustness to disorder. We conclude with a proposal that we believe will facilitate the creation of more complex states in the system under investigation and also act as a general tool to boost efficiency in QOC applications going beyond this specific system.
2 Optimization Strategy
In this work, we implement a BO training loop as illustrated in Fig. 1.
@@ 2 Optimization Strategy
Figure 1: Optimization Strategy: (top) Illustration of the Bayesian optimization training loop. The experiment or numerical simulation acts as a black-box function, providing a quality measure for a given ramping protocol. Based on evaluations of this function, the BO learner builds a surrogate model of the cost landscape and suggests the next protocol to implement. (bottom) Proposed optimization strategy exploiting numerical resources to boost optimization efficiency in the experiment. Depending on the task at hand, several avenues with varying experimental costs may be explored. The most straightforward way is to directly implement the optimization result obtained from simulation in the experiment, requiring only a single evaluation of the protocol’s quality. On top of this, the BO surrogate model can provide a sensitivity analysis that can be exploited to manually calibrate only the most sensitive protocol parameters in $\sim 10$ experimental evaluations. Finally, when substantial adaptation to the experimental setting is necessary, online optimization can be performed in $\sim 100$ to $\sim 1000$ iterations directly with the experiment. Transfer-learning approaches can leverage numerical resources to accelerate training significantly, e.g., by initializing the surrogate model used for training in the experiment with training data from the simulation. The prospective reduction of the number of cost-function evaluations in the experiment is crucial in the context of optical lattice systems where typical snapshot-based detection schemes require many experimental runs to determine an accurate figure of merit. In this work, we highlight and motivate these three scenarios by comparing the numerical simulation of an idealized system to a more realistic description of the experimental setting that takes experimental disorder into account.
@@
Gaussian processes (GP) [29] are used as a surrogate model of the cost function or figure of merit of ramping protocols over the entire parameter space. For details about BO and our implementation, the reader is referred to Appendix A.
@@ 3 Experimental Setting
Our application of BO is focused on the state-preparation problem of a bosonic two-particle FQH state recently realized in experiment [4]. The experiment features a ramp protocol, adiabatically [^1] In the context of optimization, we consider an adiabatic protocol to be any strategy aiming to minimize losses of the target-state population by limiting the ramp speed to the maximum set by the inverse of the many-body gap. transforming an initial state of two pinned particles into the delocalized FQH state.
@@
Figure 2: Experimental Ramp: (a) Experimental ramping protocol for the $4$ lattice parameters adapted from [4]. (b) Fidelity w.r.t. the instantaneous ground state throughout the ramp. The ED simulation is compared to experimental data [4]. The color-coded lines in the background show the energy levels $E-E_{0}$ of $\hat{H}(t)$ and their respective occupation during the simulated ramping protocol. Here, $E_{0}$ is the instantaneous ground-state energy. Before $t_{x}$ is turned on at $T=60\>\tau$, the system is effectively one-dimensional, since all particles are initialized in the 1st column of the $4\times 4$ lattice and there is no coupling in x-direction. In this regime, we define the fidelity wrt. this 1D system to avoid spurious drops to $F\sim 0$ when $\Delta_{x}<\Delta_{y}$ caused by the initial-state geometry.
@@
Figure 3: Optimized Ramp: (a) Optimized ramp protocol for the $4$ lattice parameters. White dots show the parameter values controlled by the optimizer at fixed points in time (vertical dashed lines). (b) Evaluated cost function at each iteration of the BO training. The visible oscillations in the evaluated fidelity follow sweeps of the BO hyperparameter $\kappa$, which tunes the strategy between exploration and exploitation (see Appendix A for optimization details). (c) Fidelity w.r.t. the instantaneous ground state throughout the ramp. The color-coded lines in the background show the energy levels of $\hat{H}(t)$ and their respective occupation during the simulated ramping protocol. All energies are w.r.t. the ground state energy which sets $E=0$. Before $t_{x}$ is turned on at $T=60\>\tau$, the system is effectively one-dimensional, since all particles are initialized in the 1st column of the $4\times 4$ lattice and there is no coupling in x-direction. In this regime, we define the fidelity wrt. this 1D system to avoid spurious drops to $F\sim 0$ when $\Delta_{x}<\Delta_{y}$ caused by the initial-state geometry.
@@
To achieve this, four global lattice parameters in the interacting Harper-Hofstadter model [30, 31, 32]
$\begin{split}\hat{H}=\ &-t_{x}\sum_{x,y}(e^{i\phi y}\hat{a}^{\dagger}_{x,y}\hat{a}^{\vphantom{\dagger}}_{x+1,y}+\mathrm{H.c.})+\Delta_{x}\sum_{x,y}x\,\hat{n}_{x,y}\\ &-t_{y}\sum_{x,y}(\hat{a}^{\dagger}_{x,y}\hat{a}^{\vphantom{\dagger}}_{x,y+1}+\mathrm{H.c.})+\Delta_{y}\sum_{x,y}y\,\hat{n}_{x,y}\\ &+\frac{U}{2}\sum_{x,y}\hat{n}_{x,y}(\hat{n}_{x,y}-1)\end{split}$ (1)
@@
are varied independently over time: (i, ii) the tunneling strengths $t_{x,y}$ in x- and y-direction and (iii, iv) the chemical potential gradients $\Delta_{x,y}$. Here, $\hat{a}^{(\dagger)}_{x,y}$ is the bosonic annihilation (creation) operator at site $(x,y)$ and $\hat{n}^{\vphantom{\dagger}}_{x,y}=\hat{a}^{\dagger}_{x,y}\hat{a}^{\vphantom{\dagger}}_{x,y}$ is the particle number operator. The tunneling phase $\phi/2\pi=0.27$ is constant throughout this work, corresponding to the Laughlin-like state [23] at the fractional filling $\nu=1/2$, where $\nu$ is the filling factor. This target state is realized in the flat lattice $\Delta_{x}=\Delta_{y}=0$, at equal tunneling amplitudes $t_{x}=t_{y}=\hbar/\tau$. We fix the value of the tunneling amplitudes in the target state as our unit of energy while the tunneling time $\tau$ acts as the unit of time. In the experiment, $\tau=4.3(1)\>\mathrm{ms}$ [4]. Finally, the on-site repulsion is fixed to $U=8\>\hbar/\tau$.
@@
The ramping protocol used in the experiment was devised by mapping out the many-body gap over the four-dimensional parameter space and manually identifying a path that avoids a gap closing. In addition, the tunneling $t_{x}$ in x-direction is only turned on in the latter half of the experimental protocol to reduce the effects of Floquet heating [33] resulting from the synthetic magnetic field engineered in the experiment. The protocol is illustrated in Fig. 2. It serves as the primary benchmark in this work – representing the capabilities of sophisticated adiabatic ramp design.
@@
where |$\ket{\psi_{\mathrm{final}}}$ is the state after the preparation protocol, and |$\ket{\psi_{0}}$ is the ground state of $\hat{H}$ at the target parameters $t_{x}=t_{y}=\hbar/\tau$ and $\Delta_{x}=\Delta_{y}=0$. This is a good figure of merit since it is experimentally accessible by inverting the ramp and measuring the fraction that has returned to the initial state using snapshots in the number basis. For the experimental protocol it was measured to be $F=43(6)\%$ [4]. In Fig. 2, we compare the experimental data to an ED time-evolution simulation. Even though the simulation is performed at zero temperature, does not take into account any lattice defects, and does not model the Floquet sequence employed in the experiment – instead simulating the unperturbed effective Hamiltonian – the fidelity-loss over time (see Fig. 2b)) shows qualitative agreement with the experiment. The fidelity of $F=53.5\%$ observed in the simulation is not much higher than the value reported in the experiment. This supports non-adiabatic losses to excited states during the state-preparation protocol as the leading source of fidelity loss, providing a promising foundation for ramp optimization.
4 Optimization
For each lattice parameter, we parametrize the time-dependent ramp by $N_{s}$ linear segments of equal length. Fixing the initial and target parameters to be the same as in the experiment leads to a BO problem with $4\times(N_{s}-1)$ parameters. Motivated by the experimental conditions and concerns about Floquet heating, we limit the search space to
@@ 4 Optimization
which corresponds to a Floquet regime that minimizes excitations to higher bands and achieves efficient dynamics. Specifically, the range for $t_{x/y}$ is chosen to limit the broadening of heating channels [33], as its strength is determined by the driving amplitude. We start ramping $t_{x}$ and $\Delta_{x}$ only after passing the time $T^{(0)}_{x}=0.6\>T$ where $T$ is the total ramp time. In summary, the optimizer may choose the values of $t_{y}$ and $\Delta_{y}$ at times $T^{(1)}_{y}=T/3$ and $T^{(2)}_{y}=2\>T/3$, and the values of $t_{x}$ and $\Delta_{x}$ at times $T^{(1)}_{x}=T^{(0)}_{x}+(T-T^{(0)}_{x})/3$ and $T^{(2)}_{x}=T^{(0)}_{x}+2\>(T-T^{(0)}_{x})/3$.
After training with $500$ iterations, BO identifies an optimal ramp with $F=94.5\%$ as illustrated in Fig. 3. Interestingly, the protocol is qualitatively similar to the one used in the experiment. As depicted in Fig. 3, it achieves near-perfect fidelity due to significant gains in the ground-state population (w.r.t. the instantaneous ground state) near the final protocol time. Such a resurgence of fidelity has also been reported in other simulation works on QOC applications [34]. While captured by optimal control theory, it is outside the scope of manual design principles, such as adiabatic ramp design, which aim to minimize fidelity losses, but do not take processes that can return population to the ground state into account. We now turn to whether this feature, observable in simulations, can be expected to be robust in experimental conditions.
5 Training with Noise
To characterize robustness, and estimate the real-world fidelity we would expect from optimized protocols, we now model experimental disorder in our simulation by the sum of random and harmonic chemical potential offsets (see Appendix C). The disorder $\mu$ is taken to be Gaussian-distributed with mean 00 and standard deviation $\sigma=0.1\>\hbar/\tau$ which is an experimentally realistic disorder level. For each numerical evaluation of a ramp’s fidelity, we perform time evolution with a newly sampled realization of the disorder.
Optimization is now performed in the presence of this disorder to represent training in an experimental setting. To mitigate fluctuations, we take the average fidelity of $5$ disorder realizations as the cost function, which is a typical value used in the literature [17, 20]. After training for $1000$ iterations, we compare the optimized protocol to the manually designed ramp and the optimum found from training without disorder in Fig. 4.